OpenStudy (anonymous):

Help Me Pleaseeeee Solve the system by using elimination or elimination with multiplication. 2x + 3y = 12 and 3x + 2y = 13

OpenStudy (cwrw238):

first multiply first equation by 3 and the second by -2 adding the 2 new equations will eliminate x solve for then plug in this value of y into the first equation to find value of x

OpenStudy (cwrw238):

* solve for y

OpenStudy (cwrw238):

can you do that gabbie?

OpenStudy (anonymous):

no lol i still dont get how to do it

OpenStudy (cwrw238):

multiplying an equation:- u need to multiply each term in the equation: 2x + 3y = 12 * 3 is 6x + 6y = 36 do you follow that?

OpenStudy (cwrw238):

i multiplied all 3 terms ( 2x , 3y and 12) by 3